OpenStudy (anonymous):

f y = x3 + 2x and dx, dt equals 5, find dy, dt when x = 2. Give only the numerical answer. For example, if dy, dt = 3, type only 3. Numerical Answers Expected!

OpenStudy (anonymous):

y = x^3 + 2 x dy/dx = 3 x^2 + 2 dy/dt = dy/dx * dx/dt = (3 x^2+2)dx/dt when dx/dt = 5 and x = 2 then dy/dt = (3*4+2)5 = 14*5 = 140/2 = 70

OpenStudy (anonymous):

thank u!!!!!!!!! :)

OpenStudy (anonymous):

You are welcome.